# A galactic outflow traced by its extended Mg II emission out to a $\sim 30$ kpc radius in the Hubble Ultra Deep Field with MUSE

### v1

@@ 8.5.2 Comparison to results from studying the CGM in absorption
Absorption lines have been the primary way to study the CGM during the last decade (e.g., results from the COS-Haloes (Werk et al., 2013) or MEGAFLOW (Schroetter et al., 2016) surveys). In particular, the MEGAFLOW survey has provided crucial understanding of the non-isotropic distribution of the CGM. Schroetter et al. (2019) use a sample of 79 strong Mg II absorbers, and find that cool halo gas exhibits a strong bimodal distribution in the azimuthal angle of the background QSO sightline, with respect to the galaxy major axis, which suggests a biconical outflow scenario that coexist with a coplanar gaseous structure. These results are further confirmed by Zabl et al. (2019). Similarly, Zabl et al. (2020) use two QSO sightlines to study the CGM of a star-forming galaxy at $z\approx 0.7$, and find that the data are consistent with an anisotropic ejection of gas from the galaxy, with a biconical outflow perpendicular to the galaxy disk.
@@ 8.5.3 Comparison to results from studying the CGM in emission
Zabl et al. (2021) report the first detection of extended Mg II emission in the halo of a galaxy that is also probed in absorption by a QSO sightline. Their analysis suggests a biconical outflow with an opening angle of $35^{\circ}$ and an outflow velocity of 130 km s-1. While this outflow velocity is almost a factor of three higher than our measured $v_{0}$, there is a fundamental difference in the assumptions made to obtain these numbers; here, we assume a radially accelerating wind, whereas in Zabl et al. (2021), the authors (motivated by the results from Bouché et al., 2012) compute this value by matching a constant velocity outflow model to the data, aiming to reproduce the absorption and emission features observed (although they also explore the possibility of radially accelerating and decelerating winds, and find that the observed surface brightness of the halo can be reproduced with either of these alternative models). Under this assumption, an outflow velocity significantly higher than our wind launching velocity is expected. Furthermore, the constant wind velocities measured by Bouché et al. (2012) from a sample of 11 Mg II absorbers are not necessarily in conflict with our results, given that the impact parameters of the QSO sightlines to the central galaxies used to obtain this result are significantly larger than the spatial extent of our outflow model. Zabl et al. (2021) also find that a Mg II halo emission powered only by scattering of continuum photons leads to an inconsistency in the density inferred from the emission and from the absorption (at an impact parameter of $\sim 40$ kpc). This result is qualitatively similar to the discrepancy we find between the observed and predicted absorption in the spectrum of BG1 (see Sec. 8.3).
@@
Burchett et al. (2021) present another study of the CGM in emission. They investigate a single galaxy at $z=0.694$, similar to #884, observed with the KCWI, that exhibits spatially resolved Mg II CGM emission extending up to a full size of $\sim 37$ kpc, with a clear P-Cygni profile in the regions where stellar continuum is present. While this work is similar to what we present in this paper (albeit with a somewhat lower spatial resolution), their conclusions differ from ours. For instance, they compare their data to a suite of 3D radiative transfer models of galactic outflows, and find that their observations are more consistent with an isotropic outflow, and radially decelerating wind. However, due to the computational expense of the radiative transfer simulations, they only explore a very limited region of the full parameter space, in terms of velocity and geometry of the wind. In addition, while in our analysis we do not explore negative values of $\gamma$ (i.e., a radially decreasing wind velocity), we find that the width of the model emission lines in the outer regions is consistent with the observed emission lines. On the other hand, Burchett et al. (2021), find that a radially increasing velocity wind would produce emission lines that are too broad compared to their observations in the outer regions, hence, a decelerating wind is required. Whether this is a particularity of the galaxy they model, or a result of exploring a limited region of the parameter space, in terms of geometry and wind velocity, has yet to be determined.
@@ 8.5.4 Comparison with results from stacked integral field data
Finally, Guo et al. (2023a) infer an outflow rate of $\dot{M}$ = $36.3\begin{subarray}{c}+16.4\\ -12.4\end{subarray}$ M yr-1, which is higher than our value reported in Sec 8.4. However, as discussed in Sec. 8.4, if the difference in the assumed dust depletion factor is accounted for, the mass outflow rate of #884 is actually higher than their measurement.
@@ 9 Summary and conclusions
    The galaxy spectrum itself is dominated by Mg II in absorption, albeit with weak P-Cygni emission humps. Towards larger impact parameters, the circumgalactic emission becomes visible and eventually dominant (Fig. 4) at surface brightnesses around and below $10^{-18}$ erg s-1 cm-2 arcsec-2 (Fig. 3), up to $\sim 30-40$ kpc from the central galaxy
@@
    We use a simple outflow modeling scheme, where we consider the outflow to be composed of thin spherical shells, whose velocity increases with radii, to fit the inner 26 kpc of the Mg II halo emission of #884 in a spatially resolved manner. We perform a MCMC approach to find the best-fitting parameters, and we produce a model cube that we compare with our data cube.
@@
    We are able to constrain the velocity profile with which the circumgalactic gas expands from the central source, described by the parameter $\gamma$, the slope of the power law dependence of velocity on radius. We find that our best model is given by $\gamma=0.7\pm 0.1$, showing that the gas velocity increases steeply towards larger distances from the central galaxy. This also demonstrates that using the Sobolev approximation to solve the radiative transfer problem is a self-consistent choice.
@@
    In terms of geometry, we find that the outflow is seen nearly face-on (i.e., the bicone pointing towards the observer, a rotation angle of $80\pm 10$), with an opening angle of $59\pm 4$. This geometry agrees well with what we would expect from a bipolar outflow, given that the central galaxy is also seen nearly face-on.
@@
    We find that the terminal velocity of the outflow is considerably higher than the escape velocity of the halo in which #884 resides, implying that the outflowing material will likely escape the gravitational potential of the galaxy, and enrich the intergalactic medium, rather than falling back onto #884. We estimate a mass outflow rate of $12\pm 7\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$, which implies a mass loading factor of $\sim 1$. However, we remind the reader that these numbers are subject to significant systematic uncertainties that come from the assumptions made to estimate them.
@@
    There are a number of differences between the data and the model. These discrepancies arise due to the difficulty for our outflow model to faithfully reproduce higher-order features of the Mg II emission, such as its complex kinematic structure in the outer regions. These differences point out that mechanisms other than outflows, such as inflows, satellites, or turbulence, could become relevant (or dominant) in shaping the CGM of galaxies, particularly towards large impact parameters.
We have shown that a relatively simple outflow modeling scheme is able to reproduce many key features of the Mg II halo around #884, and provide constraints on relevant physical mechanisms, such as the launching velocity of the galactic wind and its velocity gradient, which ultimately can be critical to understand the interplay between galaxies and their CGM. Our model reproduces the data more faithfully in the inner apertures, where the contribution from outflows is expected to be dominant. Some differences between the data and the model towards the outer apertures might indicate an increasingly relevant contribution of other phenomena shaping the CGM. The clear similarity of the circumgalactic Mg II halo around #884 to the extended Lyman-$\alpha$ envelopes of high-redshift galaxies is encouraging to use this Mg II haloes at intermediate redshifts as an analog to study and better understand the CGM emission from high-$z$ galaxies.
It would also be interesting to explore whether our simple modeling approach is also suitable for outflows with orientation angles other than nearly face-on, and in more general terms, for a more diverse set of outflows, or, alternatively, if it is worthwhile implementing a more sophisticated modeling approach, such as proper radiative transfer modeling, to faithfully reproduce the Mg II emission of a more representative sample of galaxies.
In this context, available deep VLT/MUSE data give us the opportunity to study galactic outflows in a more systematic fashion. We have already found some other galaxies that present extended Mg II emission and in a future paper we plan to extend our modeling to a larger sample of galaxies and cover a larger region of the parameter space of galaxy properties, such as stellar mass and redshift, and in this way provide statistically significant constraints to the CGM physics and the mechanisms that shape it.
@@ Appendix A Aperture-integrated spectrum modeling
In this section, we compare our results with those obtained using the same methodology described in Sec. 6, but modeling the aperture-integrated spectrum of the data cube instead of doing it in a spatially resolved manner. This is achieved simply by integrating in the spatial dimensions the data and model cubes, before calculating the likelihood, that is, computing the likelihood using only a 1D spectrum.
Figure 15 shows the posterior distribution of the model parameters obtained when we fit the aperture-integrated spectrum. Most parameters are very poorly constrained, compared to the fiducial spatially resolved modeling. Interestingly, the optical depth $\tau_{0}$ is relatively well constrained (albeit with large uncertainties), with a maximum a posteriori of $400\pm 600$. This is not unexpected, since while $\tau_{0}$ is the most impactful in the shape of a single aperture-integrated spectrum, the rest of the parameters are constrained by the spatially-resolved features of the emission, for instance, how steep is the profile, in the case of $\gamma$. However, we note that in the context of modeling the aperture-integrated spectrum, $\tau_{0}$ does not represent the optical depth at $R_{0}$, but some average quantity across the full FoV. Hence, it is significantly lower than the inferred value from our fiducial analysis.
@@ Appendix B Central aperture modeling
As discussed in Sec. 7.1, the blue wing of the Mg II absorption in the innermost aperture is particularly sensitive to the radial gradient of $\tau$, as this aperture provides a down-the-barrel view of the outflow, and changes in $\tau$ along the line of sight dominate the resultant line profile, without any additional contribution from the emission component of the P-Cygni profile. Figure 17 shows explicitly the relation between radius, velocity, flux, and optical depth, for the spectrum of the central spaxel of the model and data cubes, for the best-fitting parameters determined in Sec. 7.1.
Figure 17: Relation between radius, velocity, flux, and optical depth, for the spectrum of the central spaxel of the model and data cubes, for the best-fitting parameters determined in Sec. 7.1. Top panel shows how the optical depth $\tau$ is higher in the central region of the outflow, and decreases towards larger radii. The light blue area marks the region where radius ¡ $R_{0}$. The bottom panel shows the spectrum that results from the $\tau$ radial gradient shown above. The absorption is strongest closer to zero velocity, where the optical depth is higher, and then its strenght decreases towards more negative velocity, that correspond to larger radii and lower optical depths. Within the light blue band there is no outflow, and hence, no absorption. The figure shows the spectrum of only the Mg II $\lambda$2796 transition, and the red wing of the emission line is truncated to not include wavelength channels where the absorption of the Mg II $\lambda$2803 transition becomes relevant, making more difficult the interpretation of the figure.
@@
For comparison purposes, in this section we report the best-fitting parameters obtained when we model only the flux coming from the innermost aperture (magenta circumference in Fig. 3) of the MUSE datacube. Figure 18 shows the posterior distribution of the model parameters obtained in this case. Notably, we are still able to provide constraints in the wind properties when we only use this small portion of the datacube, that are roughly consistent with the numbers reported in our fiducial approach (see Sec. 7.1). This can be explained because although we are observing only a small region, given that the outflow is seen nearly face-on, the central sightline alone still provides information along the full outflow.
@@
The optical depth inferred in this scenario is somewhat lower than that measured in a larger FoV, and we interpret this as a consequence of the excess of emission in the outer apertures driving $\tau_{0}$ towards higher values. Due to covariance effects, a smaller $\tau_{0}$ also leads to a slightly smaller $\gamma$ and larger $v_{0}$, compared to our fiducial values. On the other hand, unsurprisingly, the geometry of the outflow is very poorly constrained from modeling only the central sightline.
@@
We have also attempted to model the full cube while keeping the value of $\tau_{0}$ fixed to the maximum posterior value obtained in this test, but this does not provide a clear improvement in the modeling of the datacube, nor in the constrain of the model parameters. Moreover, the derived parameters in this case are generally consistent with those reported in Sec. 7.1. Thus, we opt to keep $\tau_{0}$ as a free parameter in our fiducial modeling. Nevertheless, we conclude that, at least for a face-on outflow, modeling exclusively the central sightline could provide some meaningful constraints in the wind properties.
@@
This contrasts with what we report in Appendix A, that a full aperture-integrated modeling does not provide significant constraint to the outflow parameters, and might look like adding information has a negative impact on the inference of physical parameters. However, the key difference is that in the full aperture-integrated modeling, the model fitting completely loses track of where the photons come from, because photons at a given velocity (and thus, distance from the central source) could come either from a larger impact parameter but closer in the line of sight direction to the central source, or alternatively, from a smaller impact parameter, but further away in the line of sight direction, that is, distances in the plane of sky and line of sight direction cannot be disentangled. On the other hand, when modeling only the central region, different velocities can only correspond to different locations along the line of sight. Thus, paradoxically, we can track better the origin of the photons, and hence, infer the outflow properties by modeling the central sightline, provided that the outflow is (nearly) face-on. Otherwise, the central sightline would only probe a limited portion of the outflow in the velocity space.

### v3

@@ 8.5.2 Comparison to results from studying the CGM in absorption
Absorption lines have been the primary way to study the CGM during the last decade (e.g., results from the COS-Haloes (Werk et al., 2013) or MEGAFLOW (Schroetter et al., 2016) surveys). In particular, the MEGAFLOW survey has provided a crucial understanding of the non-isotropic distribution of the CGM. Schroetter et al. (2019) use a sample of 79 strong Mg II absorbers, and find that cool halo gas exhibits a strong bimodal distribution in the azimuthal angle of the background QSO sight line, with respect to the galaxy major axis, which suggests a biconical outflow scenario that coexist with a coplanar gaseous structure. These results are further confirmed by Zabl et al. (2019). Similarly, Zabl et al. (2020) use two QSO sight lines to study the CGM of a star-forming galaxy at $z\approx 0.7$, and find that the data are consistent with an anisotropic ejection of gas from the galaxy, with a biconical outflow perpendicular to the galaxy disk.
@@ 8.5.3 Comparison to results from studying the CGM in emission
Zabl et al. (2021) report the first detection of extended Mg II emission in the halo of a galaxy that is also probed in absorption by a QSO sight line. Their analysis suggests a biconical outflow with an opening angle of $35^{\circ}$ and an outflow velocity of 130 km s-1. While this outflow velocity is almost a factor of three higher than our measured $v_{0}$, there is a fundamental difference in the assumptions made to obtain these numbers; here, we assume a radially accelerating wind, whereas in Zabl et al. (2021), the authors (motivated by the results from Bouché et al., 2012) compute this value by matching a constant velocity outflow model to the data, aiming to reproduce the absorption and emission features observed (although they also explore the possibility of radially accelerating and decelerating winds, and find that the observed surface brightness of the halo can be reproduced with either of these alternative models). Under this assumption, an outflow velocity significantly higher than our wind launching velocity is expected. Furthermore, the constant wind velocities measured by Bouché et al. (2012) from a sample of 11 Mg II absorbers are not necessarily in conflict with our results, given that the impact parameters of the QSO sight lines to the central galaxies used to obtain this result are significantly larger than the spatial extent of our outflow model. Zabl et al. (2021) also find that a Mg II halo emission powered only by scattering of continuum photons leads to an inconsistency in the density inferred from the emission and from the absorption (at an impact parameter of $\sim 40$ kpc). This result is qualitatively similar to the discrepancy we find between the observed and predicted absorption in the spectrum of BG1 (see Sect. 8.3).
@@
Burchett et al. (2021) present another study of the CGM in emission. They investigate a single galaxy at $z=0.694$, similar to #884, observed with the KCWI, that exhibits spatially resolved Mg II CGM emission extending up to a full size of $\sim 37$ kpc, with a clear P-Cygni profile in the regions where stellar continuum is present. While this work is similar to what we present in this paper (albeit with a somewhat lower spatial resolution), their conclusions differ from ours. For instance, they compare their data to a suite of 3D radiative transfer models of galactic outflows, and find that their observations are more consistent with an isotropic outflow, and radially decelerating wind. However, due to the computational expense of the radiative transfer simulations, they only explore a very limited region of the full parameter space, in terms of velocity and geometry of the wind. In addition, while in our analysis we do not explore negative values of $\gamma$ (i.e., a radially decreasing wind velocity), we find that the width of the model emission lines in the outer regions is consistent with the observed emission lines. On the other hand, Burchett et al. (2021), find that a radially increasing velocity wind would produce emission lines that are too broad compared to their observations in the outer regions; hence, a decelerating wind is required. Whether this is a particularity of the galaxy they model, or a result of exploring a limited region of the parameter space, in terms of geometry and wind velocity, has yet to be determined.
@@ 8.5.4 Comparison with results from stacked integral field data
Finally, Guo et al. (2023a) infer an outflow rate of $\dot{M}$ = $36.3\begin{subarray}{c}+16.4\\ -12.4\end{subarray}$ M yr-1, which is higher than our value reported in Sec 8.4. However, as discussed in Sect. 8.4, if the difference in the assumed dust depletion factor is accounted for, the mass outflow rate of #884 is actually higher than their measurement.
@@ 9 Summary and conclusions
    The galaxy spectrum itself is dominated by Mg II in absorption, albeit with weak P-Cygni emission humps. Toward larger impact parameters, the circumgalactic emission becomes visible and eventually dominant (Fig. 4) at surface brightnesses around and below $10^{-18}$ erg s-1 cm-2 arcsec-2 (Fig. 3), up to $\sim 30-40$ kpc from the central galaxy.
@@
    We used a simple outflow modeling scheme, where we consider the outflow to be composed of thin spherical shells and whose velocity increases with radii, to fit the inner 26 kpc of the Mg II halo emission of #884 in a spatially resolved manner. We used an MCMC approach to find the best-fitting parameters, and we produced a model cube that we compared with our data cube.
@@
    We were able to constrain the velocity profile with which the circumgalactic gas expands from the central source, described by the parameter $\gamma$, the slope of the power law dependence of velocity on radius. We find that our best model is given by $\gamma=0.7\pm 0.1$, implying that the gas velocity increases steeply toward larger distances from the central galaxy. This also demonstrates that using the Sobolev approximation to solve the radiative transfer problem is a self-consistent choice.
@@
    In terms of geometry, we find that the outflow is seen nearly face-on (i.e., the bicone points toward the observer, a rotation angle of $80\pm 10$), with an opening angle of $59\pm 4$. This geometry agrees well with what we would expect from a bipolar outflow given that the central galaxy is also seen nearly face-on.
@@
    We find that the terminal velocity of the outflow is considerably higher than the escape velocity of the halo in which #884 resides, implying that the outflowing material will likely escape the gravitational potential of the galaxy and enrich the intergalactic medium, rather than falling back onto #884. We estimate a mass outflow rate of $12\pm 7\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$, which implies a mass loading factor of approximately one. However, we emphasize that these numbers are subject to significant systematic uncertainties that come from the assumptions made to estimate them.
@@
    There are a number of differences between the data and the model. These discrepancies arise due to the difficulty for our outflow model to faithfully reproduce higher-order features of the Mg II emission, such as its complex kinematic structure in the outer regions. These differences point out that mechanisms other than outflows, such as inflows, satellites, or turbulence, could become relevant (or dominant) in shaping the CGM of galaxies, particularly toward large impact parameters.
We have shown that a relatively simple outflow modeling scheme is able to reproduce many key features of the Mg II halo around #884 and provide constraints on relevant physical mechanisms, such as the launching velocity of the galactic wind and its velocity gradient, which ultimately can be critical to understanding the interplay between galaxies and their CGM. Our model reproduces the data more faithfully in the inner apertures, where the contribution from outflows is expected to be dominant. Some differences between the data and the model toward the outer apertures might indicate an increasingly relevant contribution of other phenomena shaping the CGM. The clear similarity of the circumgalactic Mg II halo around #884 to the extended Lyman-$\alpha$ envelopes of high-redshift galaxies encourages the use of these Mg II haloes at intermediate redshifts as an analog to study and better understand the CGM emission from high-$z$ galaxies.
It would also be interesting to explore whether our simple modeling approach is also suitable for outflows with orientation angles other than nearly face-on and, in more general terms, for a more diverse set of outflows. Alternatively, it would be of interest to investigate if implementing a more sophisticated modeling approach is worthwhile, such as proper radiative transfer modeling, in order to faithfully reproduce the Mg II emission of a more representative sample of galaxies.
In this context, available deep VLT/MUSE data give us the opportunity to study galactic outflows in a more systematic fashion. We have already found some other galaxies that present extended Mg II emission, and in a future paper, we plan to extend our modeling to a larger sample of galaxies and to cover a larger region of the parameter space of galaxy properties, such as stellar mass and redshift. In this way, we seek to provide statistically significant constraints to the CGM physics and the mechanisms that shape it.
@@ Appendix A Aperture-integrated spectrum modeling
In this section, we compare our results with those obtained using the same methodology described in Sect. 6, but modeling the aperture-integrated spectrum of the data cube instead of doing it in a spatially resolved manner. This is achieved simply by integrating in the spatial dimensions the data and model cubes, before calculating the likelihood, that is, computing the likelihood using only a 1D spectrum.
Figure 15 shows the posterior distribution of the model parameters obtained when we fit the aperture-integrated spectrum. Most parameters are very poorly constrained, compared to the fiducial spatially resolved modeling. Interestingly, the optical depth $\tau_{0}$ is relatively well constrained (albeit with large uncertainties), with a maximum a posteriori of $400\pm 600$. This is not unexpected, since while $\tau_{0}$ is the most impactful in the shape of a single aperture-integrated spectrum, the rest of the parameters are constrained by the spatially resolved features of the emission, for instance, how steep is the profile, in the case of $\gamma$. However, we note that in the context of modeling the aperture-integrated spectrum, $\tau_{0}$ does not represent the optical depth at $R_{0}$, but some average quantity across the full FoV. Hence, it is significantly lower than the inferred value from our fiducial analysis.
@@ Appendix B Central aperture modeling
As discussed in Sect. 7.1, the blue wing of the Mg II absorption in the innermost aperture is particularly sensitive to the radial gradient of $\tau$, as this aperture provides a down-the-barrel view of the outflow, and changes in $\tau$ along the line of sight dominate the resultant line profile, without any additional contribution from the emission component of the P-Cygni profile. Figure 17 shows explicitly the relation between radius, velocity, flux, and optical depth, for the spectrum of the central spaxel of the model and data cubes, for the best-fitting parameters determined in Sect. 7.1.
@@
For comparison purposes, in this section we report the best-fitting parameters obtained when we model only the flux coming from the innermost aperture (magenta circumference in Fig. 3) of the MUSE data cube. Figure 18 shows the posterior distribution of the model parameters obtained in this case. Notably, we are still able to provide constraints in the wind properties when we only use this small portion of the data cube, that are roughly consistent with the numbers reported in our fiducial approach (see Sect. 7.1). This can be explained because although we are observing only a small region, given that the outflow is seen nearly face-on, the central sight line alone still provides information along the full outflow.
@@
The optical depth inferred in this scenario is somewhat lower than that measured in a larger FoV, and we interpret this as a consequence of the excess of emission in the outer apertures driving $\tau_{0}$ toward higher values. Due to covariance effects, a smaller $\tau_{0}$ also leads to a slightly smaller $\gamma$ and larger $v_{0}$, compared to our fiducial values. On the other hand, unsurprisingly, the geometry of the outflow is very poorly constrained from modeling only the central sight line.
@@
We have also attempted to model the full cube while keeping the value of $\tau_{0}$ fixed to the maximum posterior value obtained in this test, but this does not provide a clear improvement in the modeling of the data cube, nor in the constrain of the model parameters. Moreover, the derived parameters in this case are generally consistent with those reported in Sect. 7.1. Thus, we opt to keep $\tau_{0}$ as a free parameter in our fiducial modeling. Nevertheless, we conclude that, at least for a face-on outflow, modeling exclusively the central sight line could provide some meaningful constraints in the wind properties.
@@
This contrasts with what we report in Appendix A, that a full aperture-integrated modeling does not provide significant constraint to the outflow parameters, and might look like adding information has a negative impact on the inference of physical parameters. However, the key difference is that in the full aperture-integrated modeling, the model fitting completely loses track of where the photons come from because photons at a given velocity (and thus distance from the central source) could come either from a larger impact parameter that is closer in the line of sight direction to the central source or from a smaller impact parameter that is farther away in the line of sight direction. In other words, distances in the plane of sky and line of sight direction cannot be disentangled. On the other hand, when modeling only the central region, different velocities can only correspond to different locations along the line of sight. Thus, paradoxically, we can track the origin of the photons better and hence infer the outflow properties by modeling the central sight line, provided that the outflow is (nearly) face-on. Otherwise, the central sight line would only probe a limited portion of the outflow in the velocity space.
Figure 17: Relation between radius, velocity, flux, and optical depth, for the spectrum of the central spaxel of the model and data cubes, for the best-fitting parameters determined in Sect. 7.1. Top panel shows how the optical depth $\tau$ is higher in the central region of the outflow, and decreases toward larger radii. The light blue area marks the region where radius ¡ $R_{0}$. The bottom panel shows the spectrum that results from the $\tau$ radial gradient shown above. The absorption is strongest closer to zero velocity, where the optical depth is higher, and then its strenght decreases toward more negative velocity, that correspond to larger radii and lower optical depths. Within the light blue band there is no outflow and hence no absorption. The figure shows the spectrum of only the Mg II $\lambda$2796 transition, and the red wing of the emission line is truncated to not include wavelength channels where the absorption of the Mg II $\lambda$2803 transition becomes relevant, making more difficult the interpretation of the figure.